OpenStudy (anonymous):

Show(1-tanx)/(1+tanx)^2 = cos2x?

OpenStudy (anonymous):

is that cos^2(x) or cos(2x)

OpenStudy (anonymous):

cos(2x)

OpenStudy (anonymous):

okay start with the trig identiy for cos(2x)

OpenStudy (anonymous):

http://prntscr.com/86l5g2 this one

OpenStudy (anonymous):

turn tan in sin/cos and play around with it, tell me if u are still stuck

OpenStudy (anonymous):

how to settle the(1+tanx)^2 i keep cant not get the answer cos2x

OpenStudy (anonymous):

To me, it is not an identity. Counterexample Let test x = pi/6, the LHS = 0.1698729811 while the RHS =0.5

OpenStudy (anonymous):

can do full step to me? i realy no idea==

OpenStudy (anonymous):

to convert (1-tanx)/(1+tanx)^2 to cos2x

OpenStudy (xapproachesinfinity):

try to simplify (1+tanx)^2 to (sec^2x+2tanx) sec^2x+2sinx/cosx

OpenStudy (xapproachesinfinity):

do like dan suggested transfer tan to sin/cos that will be helpful

OpenStudy (anonymous):

if it is (1-tan^2x)/(1+tan^2x)^2

OpenStudy (anonymous):

?

OpenStudy (xapproachesinfinity):

hmm no i was talking about the denominator of left hand side (1+tanx)^2

OpenStudy (xapproachesinfinity):

manupilate that a bit more into sin and cos

OpenStudy (xapproachesinfinity):

then after that do the same with top too

OpenStudy (anonymous):

but there is the square2 there confusing me

OpenStudy (xapproachesinfinity):

the square you need to distribute like you do with (a+b)^2

OpenStudy (anonymous):

is(1+tan^4x) equal to sec^4x?

OpenStudy (xapproachesinfinity):

no

OpenStudy (xapproachesinfinity):

there is no such identity we have 1+tan^2x=sec^2x

OpenStudy (xapproachesinfinity):

let me start a bit \(\huge \frac{1-\tan x}{\sec^2x +2\frac{\sin x}{\cos x}}=\frac{\frac{\cos x-\sin x}{\cos x}}{\frac{1}{\cos^2x}+2\frac{\sin x}{\cos x }}\)

OpenStudy (xapproachesinfinity):

see if you can finish it off

OpenStudy (xapproachesinfinity):

you can cancel that cos that is in top and bottom

OpenStudy (irishboy123):